OpenStudy (jfruihvifjokp):

Point G(−7, 4)G(−7, 4) is translated using the rule (x+10, y−6)(x+10, y−6). What is the x-coordinate of G′G′ ? Enter your answer in the box.

OpenStudy (markchernioglo):

(x+7)(y+4) (x+7)(x+4)

OpenStudy (markchernioglo):

Just figured it out. :)

OpenStudy (jfruihvifjokp):

wait so whats the answer

OpenStudy (markchernioglo):

it is what I first posted above. :)

OpenStudy (jfruihvifjokp):

all tht dont fit

OpenStudy (markchernioglo):

Ok ok so I think they want it this way then, (x+7,y+4)

OpenStudy (markchernioglo):

Let me know if that works.

OpenStudy (jfruihvifjokp):

nope

OpenStudy (markchernioglo):

Huh, its supposed to be correct.. @TheSmartOne can you check?

TheSmartOne (thesmartone):

The rule is (x+10, y−6) The coordinate of G is (-7, 4) What is -7 + 10 = ? What is 4 - 6 = ?

OpenStudy (markchernioglo):

Oooooh, I see now.

OpenStudy (jfruihvifjokp):

it was actually 1,3

TheSmartOne (thesmartone):

It should be (3, -2)

OpenStudy (markchernioglo):

Yea, because -7+10=3 and 4-6=-2

OpenStudy (jfruihvifjokp):

idek wen i checked it said 1,3